# NotebookOS: A Notebook Operating System for Interactive Training with On-Demand GPUs

### v1

@@ 5.1.2. Setup
We conducted our evaluation on AWS EC2 (ec2, ). We ran each baseline on a cluster of 30 GPU EC2 VMs. To be consistent with the Sensei production setups, each VM is equipped with 8 GPUs. We implemented a workload driver and associated dashboard to streamline the deployment and execution of workloads on NotebookOS. To ensure a thorough evaluation, we followed (elasticflow_asplos23, ) and integrated support for a variety of different DL models and datasets across several different application domains (Table 1). The workload driver randomly assigns each client an application domain, after which a random dataset and model are assigned. Each cell task request submitted by the client will train its assigned model on the assigned dataset, which are retrieved by kernel replicas from an AWS S3 bucket.
We evaluate NotebookOS using a 17.5-hour excerpt of the Sensei workload trace. To further elucidate its performance and monetary cost implications of NotebookOS, we also implemented a robust, detailed simulator that we used to run the full Sensei trace. The simulator implemented NotebookOS’s default scheduling/placement policy along with the baseline policies described in section 5.1.1. The results of our simulation study on NotebookOS are presented in section 5.5.
@@ 5.3. Active Sessions & Training Events
Figure 7. The number of active user-submitted training tasks and active user sessions during the 17.5-hour SenseiTrace.
Figure 7 presents a timeline plot showing the number of active sessions during the execution of the 17.5-hour Sensei workload trace excerpt on NotebookOS. This series is plotted on the secondary (i.e., right) y-axis. The number of active sessions increases from 0 at the beginning of the trace excerpt to 87 by the end of the excerpt. The maximum number of active sessions at any given time throughout the 17.5-hour SenseiTrace excerpt is 90.
Figure 7 also presents a timeline plot showing the number of user-submitted training events being processed during the execution of the 17.5-hour SenseiTrace excerpt on NotebookOS. This series is plotted on the primary (i.e., left) y-axis. Initially there are no active, user-submitted trainings. At the end of the trace excerpt, there are 26 active user-submitted trainings. The mean and medium number of active, user-submitted trainings are 19.5 and 19, respectively. The maximum number of active user-submitted trainings at any given time throughout the summer SenseiTrace is 141 (see Figure 19 in appendix A). The maximum number of active user-submitted trainings at any given time throughout the 17.5-hour SenseiTrace excerpt is 34.
Figure 8. GPU usage timelines. Curves for Batch (left), NotebookOS (middle), and NotebookOS (LCP) (right) represent the number of GPUs provisioned under each policy, while “oracle” represents an optimal policy, which provisions the exact number of GPUs required to serve training requests. The number of GPUs saved relative to Reservation is shown in the green-shaded region, while the number of GPUs over-provisioned compare to “oracle” are represented by the orange-shaded region.
@@
Figure 9. CDFs of (a) interactivity delays and (b) task completion times (TCT) across different scheduling policies.
@@ 5.3.1. Interactivity
The Reservation baseline binds GPUs exclusively to the user’s notebook kernel for its entire lifetime, therefore it provides a high degree of interactivity for users. However, these resources are typically underutilized during IDLT workloads, as users often spend long intervals debugging, modifying code, or stepping away while GPUs remain idle but reserved.
Effect of Multiple Kernel Replicas. NotebookOS commits GPUs to a kernel replica immediately upon receiving an execution request 89.6% of the time. Similarly, 89.45% of the time, NotebookOS reused the same executor replica for consecutive “execute_request” messages. Using its replicated kernel design, NotebookOS maximizes the chances that at least one replica has available resources when a training request arrives, achieving nearly the same interactivity as the reservation baseline.
@@ 5.3.2. Task Completion Time (TCT)
TCT measures the interval between the instant a cell task is submitted to the instant the kernel completes the task execution. As shown in Figure 9(b), NotebookOS achieves a TCT distribution comparable to Reservation. NotebookOS has slightly higher TCT from the $38^{th}$ to the $90^{th}$ percentile. This is because some sessions were not able to find available GPUs on their servers (due to oversubscription and server’s full GPU usage) and they are migrated to different servers. When NotebookOS’s pre-warmed containers were exhausted, container provisioning were created on demand, incurring long cold startup delays. This can also be observed from the tail of NotebookOS’s interactivity delays in Figure 9(a). NotebookOS (LCP) sees much longer TCTs as a submitted cell request triggered a warming-up operation to download model parameters and datasets. Batch results in the highest TCTs because kernel replica containers are provisioned on demand for every request, and model parameters and datasets must be read and written before and after processing each code submission.
5.3.3. Resource Usage & Efficiency
Figure 10. Timeline of major events occurred while executing the 17.5-hour workload on NotebookOS, with NotebookOS’s subscription ratio plotted on the secondary Y-axis.
Batch (Figure 8 left), achieves significantly improved resource utilization compared to Reservation because Batch only provisions containers and allocates resources in response to training requests. NotebookOS achieves a high GPU utilization compared to Reservation (Figure 8 middle). Specifically, NotebookOS and NotebookOS (LCP) saved a total of 1,187.66 and 1,662.53 GPU hours, respectively. However, because NotebookOS creates 3 long-running kernel replicas and maintains a small buffer of “extra” GPU servers to accommodate bursts of training requests, NotebookOS ultimately provisions more servers than Batch. Specifically, NotebookOS provisions kernel replicas as long as the associated sessions are alive, while Batch only provisions kernels while those sessions are training. NotebookOS (LCP) (Figure 8 right) enhances NotebookOS’s resource efficiency by provisioning roughly 23.52% fewer GPUs by the end of the workload, though it still provisions 18.18% more GPUs than Batch.
@@ 5.3.4. Subscription Ratio
The relationship between the cluster-wide subscription ratio (SR, described in section 3.4) and the triggering of scale-out events by NotebookOS’s auto-scaling policy (section 3.4.2) are clearly shown in Figure 10. Specifically, the SR increases sharply at the beginning of the workload when many new Distributed Kernels are created. To accommodate these kernel replicas, NotebookOS triggers scale-out events. As new servers are provisioned, the cluster-wide subscription ratio falls. Around hour two, there is another burst of kernel replica creations, leading to another increase in the SR. Subsequently, additional servers are provisioned to serve the new kernel replicas shortly after the fourth hour, leading to another drop in the SR. A similar pattern occurs shortly before the twelfth hour. Similarly, there is an increase in the frequency of kernel migrations when the SR begins to climb, reflecting growing GPU resource saturation and potential contention when training requests arrive. Notably, a spike in migrations coincides with the first spike in the SR, followed by additional migrations after the fourth and twelfth hours of workload execution.
@@ 5.4. Object Synchronization Overhead
Figure 11. CDF of the latency of reading and writing large objects and synchronizing smaller object state via the Raft synchronization protocol. Note that the x-axis is log-scale.
Figure 11 presents a timeline plot showing the number of active sessions during the execution on NotebookOS of the 17.5-hour SenseiTrace excerpt. The “Sync” curve corresponds to the end-to-end latency of synchronizing small objects via the Raft protocol. (This process is described in detail in section 3.2.4.) The “Sync” latency is typically extremely short: the $90^{th}$, $95^{th}$, and $99^{th}$ percentile values are 54.79ms, 66.69ms, and 268.25ms, respectively. The “Reads” and “Writes” curves correspond to the latencies observed when reading and writing large objects from and to the Distributed Data Store, respectively. These operations are reasonably short, as 99% of read and write operations complete within roughly 3.95 and 7.07 seconds, respectively.
As discussed in section 2.3, the $50^{th}$ and $75^{th}$ percentile of task IATs are 300 seconds (5 minutes) and 480 seconds (8 minutes) for the SenseiTrace, respectively. The shortest event IAT within the SenseiTrace is 240 seconds, or 4 minutes. Notably, the overhead of writing and reading large objects to and from the Distributed Data Store is completely contained within event IATs. As a result, this overhead is almost completely hidden from users in IDLT workloads such as the Sensei workload.
@@ 5.5.1. Monetary Cost
To understand the cost implications of NotebookOS, we compare it with Reservation in terms of provider costs—specifically, the operational cost of provisioning EC2 resources, as done by production notebook platforms like Sensei)—and the revenue generated by the notebook service provider.
We implemented a simple billing model, assuming the provider covers AWS EC2 VM costs. Users pay 1.15× the provider’s rate, proportional to resource usage. Standby Distributed Kernel replicas are charged 12.5% of the base rate. For example, if a provider pays \$10/hour for an 8-GPU EC2 VM, each standby replica is billed \$1.44/hour ($10\times 1.15\times 0.125$). When a replica runs a training task using 4 GPUs, the charge increases to \$5.75/hour ($10\times 1.15\times 0.5$). Reservation follows the same 1.15$\times$ multiplier for on-demand GPU usage.
@@
Figure 12. Timeline of provider cost (provisioned EC2 resources), revenue, and profit margin.
@@ 6. Related Work
Notebook Platforms. Two types of notebook platforms emerge, reservation-based and batch-based. Commercial notebooks like Google Colab (colaboratory, ), NaaS (naas, ), Lentiq (lentiq, ), and JupyterHub (jupyterhub, ) offer flexible GPU reservations but suffer from low utilization. To mitigate this, they use keep-alive to terminate inactive sessions, which helps reclaim resources but does not resolve the core utilization issue. Batch-connect notebook services, such as Azure Machine Learning (azure_ml, ) and others (ncar_papermill, ; sagemaker_notebooks, ; sagemaker_studio_lab, ), submit GPU jobs via batch schedulers. However, this incurs cold startup overhead and scheduling delays, disrupting interactive programming.
GPU Cluster Schedulers.
@@
State-of-the-art batch GPU cluster schedulers (gandiva_osdi18, ; tiresias_nsdi19, ; themis_nsdi20, ; gavel_osdi20, ; pollux_osdi21, ) aim to balance the competing demands of long-running GPU training jobs but overlook responsiveness and interactivity, making them poorly suited for notebook IDLT workloads.
@@
Serverless. Optimized snapshot-loading reduces cold start costs for CPU serverless functions (sock_atc18, ; faasnap_eurosys22, ; vhive_asplos21, ; aws_lambda_snapstart, ) and serverless GPUs (parallel_gpu_os, ; serverlessllm_osdi24, ; lambdascale_arxiv25, ). NotebookOS’s pre-warmed container pool can leverage these methods to further cut GPU startup cost.
@@
GPU Virtualization. GPU virtualization solutions use NVIDIA MIG (nvidia_mig, ), MPS (nvidia_mps, ), and CUDA API interposition for performance (tgs_nsdi23, ; tally_asplos25, ), memory isolation (faaswap, ), and parallel execution on shared GPUs (rcuda_hpcs10, ). Currently, NotebookOS supports only coarse-grained, exclusive GPU time-sharing. Fine-grained GPU sharing is part of our future work.

### v2

@@ 5.1.2. Setup
We conducted our evaluation on AWS EC2 (ec2, ). We ran each baseline on a cluster of 30 GPU EC2 VMs, each equipped with 8 GPUs to match the Adobe research cluster production setups. We developed a workload driver and dashboard to automate workload deployment and execution on NotebookOS. Following (elasticflow_asplos23, ), we integrated support for a diverse set of DL models and datasets across several different application domains (Table 1). The workload driver randomly assigns each client an application domain, after which a random dataset and model are assigned. Each cell task request submitted by the client will train its assigned model on the assigned dataset, which are retrieved by kernel replicas from an AWS S3 bucket.
We evaluate NotebookOS using a 17.5-hour excerpt of the Adobe research cluster workload trace. To further elucidate its performance and monetary cost implications of NotebookOS, we also implemented a robust, detailed simulator that we used to run the full Adobe research cluster trace. The simulator implemented NotebookOS’s default scheduling/placement policy along with the baseline policies described in section˜5.1.1. The results of our simulation study on NotebookOS are presented in section˜5.5.
@@ 5.3. Active Sessions & Training Events
Figure 7. The number of active user-submitted training tasks and active user sessions during the 17.5-hour AdobeTrace.: The number of active user-submitted training tasks and active user sessions during the 17.5-hour AdobeTrace.
Figure 8. Provisioned GPUs timelines. Curves for Batch (left), NotebookOS (middle), and NotebookOS (LCP) (right) represent the number of GPUs provisioned under each policy, while “oracle” represents an optimal policy that provisions the exact number of GPUs required to serve training requests. The GPUs saved relative to Reservation is shown in the green-shaded region, while the number of GPUs over-provisioned compared to “oracle” are represented by the orange-shaded region.: GPU usage timelines.
Figure 7 presents a timeline plot showing the number of active sessions during the execution of the 17.5-hour workload trace excerpt on NotebookOS. This series is plotted on the secondary (i.e., right) y-axis. The number of active sessions increases from 0 at the beginning of the trace excerpt to 87 by the end of the excerpt. The maximum number of active sessions at any given time throughout the 17.5-hour AdobeTrace excerpt is 90.
Figure 7 also presents a timeline plot showing the number of user-submitted training events being processed during the execution of the 17.5-hour AdobeTrace excerpt on NotebookOS. This series is plotted on the primary (i.e., left) y-axis. Initially there are no active, user-submitted trainings. At the end of the trace excerpt, there are 26 active user-submitted trainings. The mean and medium number of active, user-submitted trainings are 19.5 and 19, respectively. The maximum number of active user-submitted trainings at any given time throughout the summer AdobeTrace is 141 (see Figure 20 in appendix˜A). The maximum number of active user-submitted trainings at any given time throughout the 17.5-hour AdobeTrace excerpt is 34.
5.3.1. Resource Usage & Efficiency
Batch (Figure 8 left), achieves significantly improved resource utilization compared to Reservation because Batch only provisions containers and allocates resources in response to training requests. NotebookOS achieves higher GPU utilization than Reservation (Figure 8, middle), saving 1,187.66 (NotebookOS) and 1,662.53 (NotebookOS (LCP)) GPU hours. However, because NotebookOS maintains three long-running replicas per kernel as well as a small buffer of “extra” GPU servers for request bursts, NotebookOS provisions more servers than Batch, which only allocates GPUs during cell executions. NotebookOS (LCP) (Figure 8, right) improves resource efficiency, provisioning 23.52% fewer GPUs than NotebookOS but still 18.18% more than Batch. Note that the oracle curve shows the number of GPUs required to serve all active training requests, rather than aggregated GPU bandwidth utilization, which is shown in Figure 2(d). NotebookOS provisions slightly more GPUs than Batch/LCP to account for its use of replicated kernels and to handle request bursts.
@@ 5.3.1. Resource Usage & Efficiency
Figure 9. CDFs of (a) interactivity delays and (b) task completion times (TCT) across different scheduling policies.: CDFs of (a) interactivity delays and (b) task completion times (TCT) across different scheduling policies.
@@ 5.3.2. Interactivity
The Reservation baseline binds GPUs to a user’s notebook kernel for its full lifetime, ensuring high interactivity. However, IDLT workloads often under-utilize GPU resources, as users frequently debug, edit code, or pause while GPUs remain idle but reserved.
Effect of Multiple Kernel Replicas. NotebookOS commits GPUs to a kernel replica immediately upon receiving an execution request 89.6% of the time. Similarly, 89.45% of the time, NotebookOS reused the same executor replica for consecutive execution request. NotebookOS’s use of replicated kernels maximizes the chances that at least one replica has available resources when a training request arrives, achieving nearly the same interactivity as the reservation baseline.
@@ 5.3.3. Task Completion Time (TCT)
TCT measures the interval between cell submission and the instant the cell’s execution is completed. As shown in Figure 9(b), NotebookOS achieves a TCT distribution comparable to Reservation, with slightly higher TCTs from the $38^{th}$ to the $90^{th}$ percentile. This is because some sessions were not able to find available GPUs on their servers (due to oversubscription and server’s full GPU usage) and thus are migrated to different servers. When NotebookOS’s pre-warmed containers were exhausted, new containers were created on demand, incurring long cold startup delays. This can also be observed from the tail of NotebookOS’s interactivity delays in Figure 9(a). NotebookOS (LCP) sees much longer TCTs as a submitted cell request triggered a warming-up operation to download model parameters and datasets. FCFS yields the highest TCTs due to on-demand notebook kernel provisioning and mandatory pre- and post-processing data I/O (model parameters and datasets) for each submission.
@@ 5.3.4. Subscription Ratio
Figure 10. Timeline of major events occurred while executing the 17.5-hour workload on NotebookOS, with NotebookOS’s subscription ratio plotted on the secondary Y-axis.: Timeline of major events occurred while executing the 17.5-hour workload on NotebookOS, with NotebookOS’s subscription ratio plotted on the secondary Y-axis.
The relationship between the cluster-wide subscription ratio (SR, described in section˜3.4) and the triggering of scale-out events by NotebookOS’s auto-scaling policy (section˜3.4.2) are clearly shown in Figure 10. Specifically, the SR increases sharply at the beginning of the workload when many new Distributed Kernels are created. To accommodate these kernel replicas, NotebookOS triggers scale-out events. As new servers are provisioned, the cluster-wide SR falls. Around hour two, there is another burst of kernel replica creations, leading to another increase in the SR. Subsequently, additional servers are provisioned to serve the new kernel replicas shortly after hour four, leading to another drop in the SR. A similar pattern occurs before hour twelve. Similarly, there is an increase in the frequency of kernel migrations when the SR begins to climb, reflecting growing GPU resource saturation and potential contention when training requests arrive. Notably, a spike in migrations coincides with the first spike in the SR, followed by additional migrations after hours four and twelve of workload execution.
@@ 5.4. Object Synchronization Overhead
Figure 11. CDF of the latency of reading and writing large objects and synchronizing smaller object state via the Raft synchronization protocol. Note that the x-axis is log-scale.: CDF of the latency of reading and writing large objects and synchronizing smaller object state via the Raft synchronization protocol. Note that the x-axis is log-scale.
Figure 11 presents a timeline plot showing the number of active sessions during the execution on NotebookOS of the 17.5-hour AdobeTrace excerpt. The “Sync” curve corresponds to the end-to-end latency of synchronizing small objects via the Raft protocol (section˜3.2.4). The “Sync” latency is typically extremely short: the $90^{th}$, $95^{th}$, and $99^{th}$ percentile values are 54.79ms, 66.69ms, and 268.25ms, respectively. The “Reads” and “Writes” curves correspond to the latencies observed when reading and writing large objects from and to the Distributed Data Store, respectively. These operations are reasonably short, as 99% of read and write operations complete within roughly 3.95 and 7.07 seconds, respectively.
As discussed in section˜2.3, the $50^{th}$ and $75^{th}$ percentile of task IATs are 300 seconds (5 minutes) and 480 seconds (8 minutes) for the AdobeTrace, respectively. The shortest event IAT within the AdobeTrace is 240 seconds, or 4 minutes. Notably, the overhead of writing and reading large objects to and from the Distributed Data Store is completely contained within event IATs. As a result, this overhead is almost completely hidden from users in IDLT workloads such as the Adobe research cluster workload.
@@ 5.5.1. Monetary Cost
To understand the cost implications of NotebookOS, we compare it with Reservation in terms of provider costs—specifically, the operational cost of provisioning EC2 resources, as done by production notebook platforms like Adobe research cluster—and the revenue generated by the notebook service provider.
[…]
@@
Figure 12. Timeline of provider cost (provisioned EC2 resources), revenue, and profit margin.: Timeline of provider cost (provisioned EC2 resources), revenue, and profit margin.
5.5.2. GPU Hours Saved
Figure 13. The number of GPU-hours saved by NotebookOS by avoiding unnecessary re-execution of notebook cells following idle reclamations of user sessions.: The number of GPU hours saved by NotebookOS by avoiding unnecessary re-execution of notebook cells following idle reclamations of user sessions.
Figure 13 illustrates the number of GPU hours saved by NotebookOS by avoiding unnecessary re-execution of notebook cells following idle session reclamations across the entire AdobeTrace. Each of the five lines in the plot represents a different configuration of the idle reclamation interval—the duration for which a Jupyter Notebook session may remain idle before its associated kernel is reclaimed. In the absence of NotebookOS’s state replication and persistence mechanisms, reclaiming idle sessions would result in the loss of in-memory state, requiring cell re-execution and additional GPU usage when users return. By preserving session state, NotebookOS prevents this redundant computation and significantly reduces GPU waste.
@@ 6. Related Work
Notebook Platforms.
Commercial notebook services such as Google Colab (colaboratory, ), NaaS (naas, ), Lentiq (lentiq, ), and JupyterHub (jupyterhub, ) allow users to reserve GPUs for however long they like with great flexibility but low GPU utilization. These platforms add keep-alive and timeout strategies so that the platform can automatically terminate a notebook instance if the user does not interact with it for a configurable period of time. The keep-alive strategy can help save unused GPU resources to some extent but does not fundamentally address the low GPU utilization issue. Azure Machine Learning (azure_ml, ), Open OnDemand (open_ondemand, ; osc_jupyter_batchconnect, ), among others (ncar_papermill, ; sagemaker_notebooks, ; sagemaker_studio_lab, ), offer “batch-connect” notebook services to allow users to submit GPU jobs through a batch scheduler. This approach introduces significant runtime cold startup cost and batch scheduling delays during request bursts, undermining the interactive nature of notebook programming.
@@
GPU Cluster Schedulers. State-of-the-art batch GPU cluster schedulers (gandiva_osdi18, ; tiresias_nsdi19, ; themis_nsdi20, ; gavel_osdi20, ; pollux_osdi21, ) aim to balance the competing demands of long-running GPU training jobs but overlook responsiveness and interactivity, making them poorly suited for notebook IDLT workloads. Gandiva (gandiva_osdi18, ) is a GPU cluster scheduler that leverages domain-specific knowledge and an introspective scheduler to dynamically optimize BDLT workloads. Tiresias (tiresias_nsdi19, ) is a GPU cluster manager that minimizes job completion time by prioritizing jobs with partial information and tasks with the least service received. Themis (themis_nsdi20, ) employs a 2-tiered, “finish-time fair” approach where workloads bid for GPUs in an auction, balancing short-term efficiency with long-term fairness.
@@
Serverless. Optimized snapshot-loading reduces cold start costs for CPU (sock_atc18, ; faasnap_eurosys22, ; vhive_asplos21, ; aws_lambda_snapstart, ) and GPU (parallel_gpu_os, ; serverlessllm_osdi24, ; lambdascale_arxiv25, ) serverless functions. Locality-driven keep-alive (peeking_atc18, ; faascache_asplos21, ; rainbowcake_asplos24, ; cidre_asplos25, ) optimize container caching for minimizing cold start costs. NotebookOS’s pre-warmed container pool can leverage these methods to further cut GPU startup cost.
@@
GPU Virtualization. Existing solutions for GPU virtualization and sharing leverage NVIDIA MIG (nvidia_mig, ) / MPS (nvidia_mps, ) as well as user-space CUDA API interposition and remoting techniques for GPU kernel performance isolation (tgs_nsdi23, ; tally_asplos25, ), memory isolation (faaswap, ), and parallel executions within a single remote GPU (rcuda_hpcs10, ). Currently, NotebookOS supports only coarse-grained, exclusive GPU time-sharing via kernel replication and executor replica election but does not support fine-grained, concurrent, spatial GPU sharing. Enabling fine-grained GPU sharing is part of our future work.
Consensus Protocols and Leader Election Algorithms. Raft (raft_atc14, ) is a widely-adopted consensus algorithm designed for understandability while providing strong consistency, featuring leader elections with randomized timeouts (to prevent ties) and log replication via heartbeats. NotebookOS’s state synchronization and executor selection protocols are implemented atop Raft, as they require more complex logic than what is provided directly by Raft’s leader elections. The use of Raft provides robust, off-the-shelf consistency and fault tolerance support. Alternative consensus protocols, such as Viewstamped Replication (viewstamped_podc88, ), EPaxos (epaxos_sosp13, ), and Zab (zab_dsn11, ), offer tradeoffs in latency, leader dependence, and throughput. The Bully algorithm (bully_tc82, ) targets crash-recovery systems and elects the node with the highest ID, while the Ring algorithm (ring_election_algo, ; ring_algo_lelann, ) organizes nodes into a logical ring in which communication occurs only between adjacent neighbors. Systems like ZooKeeper (zookeeper_atc10, ) and Chubby (chubby_osdi06, ) offer practical leader election mechanisms, informing the design of resilient control planes in NotebookOS.
7. Limitations and Future Work
NotebookOS does not yet support GPU sharing or fractional allocations. Future work will incorporate GPU virtualization using tools such as HAMi (hami, ). While multi-server training is also not supported, it can be enabled through integration with distributed training frameworks like Ray (ray_osdi18, ). These capabilities are largely orthogonal to NotebookOS’s core abstractions, and NotebookOS’s mechanisms for scheduling and execution would remain largely unchanged. Lastly, NotebookOS could further optimize CPU-only notebook cells by enabling elastic burst-parallel executions through the integration of stateful serverless scheduling techniques (wukong_socc20, ; cloudburst_vldb20, ).